OpenStudy (anonymous):

When a block is placed on top of a vertical spring, the spring compresses 4.40 cm. Find the mass of the block, given that the force constant of the spring is 1700 N/m.

OpenStudy (anonymous):

In a position of equilibrium , The spring force is equal to the weight So, kx = mg m=kx/g = (1700)(4.4 /100)/ 9.8

OpenStudy (anonymous):

This is a question on spring constant. So F = kx mg=kx m=kx/g =(1700)(0.044)/(100) = NB: Change 4.40cm to m so you'll divide 4.40 over 100.

OpenStudy (anonymous):

1700*(0.044)=M*(9.8) so M=7.63kg

OpenStudy (anonymous):

at equillibrium force applied by spring will be equal to weight of the block. so, kx=mg, and given k=1700N/M, x=.044M 1700*.044 = weight of the block, 9.8*m = 1700*.044 where m is the mass of the block.